40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 7 03:14:06 2013, Jenne, Update, SUS, PRM SUS_LSC violin (FM5) set to correct frequency 
    Reply  Wed Aug 21 23:50:40 2013, Koji, Update, SUS, PRM SUS_LSC violin (FM5) set to correct frequency 
       Reply  Fri Aug 23 01:52:34 2013, Jenne, Update, SUS, Violin filters moved to LSC, from SUS MICHloopNotches_23Aug2013.pdf
          Reply  Fri Aug 23 14:19:28 2013, rana, Update, SUS, Violin filters moved to LSC, from SUS 
Message ID: 9057     Entry time: Fri Aug 23 01:52:34 2013     In reply to: 9048     Reply to this: 9058
Author: Jenne 
Type: Update 
Category: SUS 
Subject: Violin filters moved to LSC, from SUS 

[Rana, Jenne]

We were meditating a little bit on what may be the story behind the PRM violin filter situation.  We locked the PRMI, and turned on and off the violin filters.  We noticed, very bizarrely, that when the violin filters were ON, the servos would oscillate.  Weird.  Also, probably because the oscillation was causing us to hit the limit we have in the MICH servo, we rung up a 3rd harmonic of one of the violin modes, which was at 1955 Hz. 

We took a transfer function of the PRCL servo, saw that the UGF was 300 Hz, and lowered it to ~180 Hz.  After later investigations, that high-ish UGF probably wasn't a problem.  Anyhow, we then took MICH servo transfer functions, and saw some very weird stuff. 

At frequencies where we had violin filter notches, we were seeing peaks in the transfer function, which came close to touching, or crossed the 0dB line!  We suspect that this may have something to do with the balancing of the drives to the optics, since we have PRCL driving PRM, but MICH driving BS and PRM.  What we did was move the violin filter notches into the LSC model.  There were already SUS filter banks in the LSC model (right side of the LSC screen).  In preparation for the DRMI, I have put the BS violin notches into the BS, PRM and SRM filter banks, as well as the PRM and SRM filters into all 3 banks.  Right now for PRMI, I have the BS and PRM notches (as well as the Vio3 notch) turned on in both BS and PRM.  All of the violin-related filters are turned off in the LSC filter bank inside the SUS models.  When we did this, the servo oscillations no longer are excited when we turn on the notches, and when we take a new transfer function, there are no longer weird peaks at the notch frequencies.  More meditation needs to be done.

Also, for the violin3 filter, Rana noted that at 1955 Hz, after we confirmed that the REFL 55 phase was set correctly (and we're using REFL 55 I&Q for PRMI locking), the I-phase had a response of 0.36, while the Q-phase had a response of 0.20.  I should be able to think about these numbers, and decide if the vio3 is for the BS or the PRM violin mode.

 MICHloopNotches_23Aug2013.pdf

the above series of Bode plots shows the MICH Open Loop Gain as the REFL55 phase is adjusted (PURPLE, ORANGE) with the notches in the SUS and then (RED) as the notches are moved to the LSC and made the same for all optics.

In other news, I have the parts that Jamie ordered to make a new 110 demod board, so that's one of my daytime activities now, so we can have both POP110 and AS110.

ELOG V3.1.3-